OpenStudy (anonymous):

Which of the following is a quadratic function? A)y = x2 – 20 B)y = x3 – 25x C) y = x4 + 13x D)y = x – 49

OpenStudy (anonymous):

Did you mean \[y = x^2 -20\]\[y = x^3 -25x\]\[y = x^4 +13x\]\[y = x-49\]

OpenStudy (anonymous):

It's okay. In pellet case have look at the highest exponent in each term and look if it's squared.

OpenStudy (anonymous):

So for example \(y = x^3-25x\) is cubic.

OpenStudy (anonymous):

how is that a cubic

OpenStudy (anonymous):

Because the highes exponent is \(3\)

OpenStudy (anonymous):

is it C) y = x^4 + 13x

OpenStudy (anonymous):

Okay lets first make yourself clear what a quadratic function is.

OpenStudy (anonymous):

So \(x^1\) is a linear function, \(x^2\) is called quadratic and \(x^3\) is cubic